OpenStudy (anonymous):

please help? Evaluate the definite integral (if it exists) intergral (e^1/3)/-6x^2 dx from 1 to 2 ?

OpenStudy (amistre64):

e^1/3 looks like a variable, but its not; its just a constant

OpenStudy (anonymous):

\[\int\limits\limits_{1}^{2} \frac{ e^\frac{ 1 }{ x } }{ -6x^2 } dx\]

OpenStudy (amistre64):

\[\huge \int\limits\limits_{1}^{2} \frac{ e^\frac{ 1 }{ x } }{ -6x^2 } dx\] is that 1/3 or 1/x? need to verify

OpenStudy (anonymous):

thats what it looks like properly :)

OpenStudy (anonymous):

\[\frac{ 1 }{ x }\]

OpenStudy (amistre64):

well, if this is going to be simple than with any luck it will come from something like:\[\Large \frac d{dx}e^{1/x}\]

OpenStudy (amistre64):

so lets start by taking the derivative of that

hartnn (hartnn):

yeah, just put u= 1/x du=... ?

OpenStudy (anonymous):

exactly, @hartnn, this substution does the trick.

OpenStudy (anonymous):

do we not have to do something with the derivavtive of -6x^2?

OpenStudy (amistre64):

we should verify that if its going to be simple first and see if we can modifiy it with a useful form of "1"

OpenStudy (anonymous):

the integral exists anyway! i know that much

OpenStudy (amistre64):

\[\frac d{dx}e^{1/x}=\frac{e^{1/x}}{-2x^2}\]agree?

OpenStudy (anonymous):

where did the -2x^2 come from? is the derivative of -6x^2 not 12x?

OpenStudy (amistre64):

your confusing your rules .... its best to see these things on a more holistic level :)

OpenStudy (anonymous):

oh ok :)

OpenStudy (amistre64):

\[\frac d{dx}e^{u(x)}=\frac{d}{dx}u(x)~e^{u(x)}=u' ~e^u\]

OpenStudy (amistre64):

let u= 1/x du/dx = -1/2x^2 du = -1/2x^2 dx

OpenStudy (anonymous):

how do i get the derivative of 1/x?

OpenStudy (amistre64):

1/x = x^-1 which is then just the power rule .... which means i got a bad 2 in there :)

OpenStudy (amistre64):

im thinking of a sqrt function ..... silly me

OpenStudy (anonymous):

what do you mean a bad 2? :)

OpenStudy (amistre64):

x^-1 goes to -x^(-2) 1/x goes to -1/x^2 NOT -1/2x^2

OpenStudy (anonymous):

oh ok i get ya now

OpenStudy (amistre64):

so just pull out the 1/6 and integrate the easy way

OpenStudy (anonymous):

wheres the 1/6?

OpenStudy (amistre64):

.....

OpenStudy (amistre64):

\[\int\frac{e^{1/x}}{-6x^2}dx\] \[\int\frac16 \frac{e^{1/x}}{-x^2}dx\] \[\frac16\int \frac{e^{1/x}}{-x^2}dx\]

OpenStudy (anonymous):

oh yes i see :) got confused for a minute!

OpenStudy (amistre64):

just becasue you were doing vector calculus the other day doesnt mean you have to forget the fundamentals ;)

OpenStudy (amistre64):

you got it from here?

OpenStudy (anonymous):

\[\frac{ e^\frac{ 1 }{ x } }{ 6 }\]

OpenStudy (anonymous):

something like that right?

OpenStudy (amistre64):

yes, but apply your limits

OpenStudy (anonymous):

what do you mean?

hartnn (hartnn):

e^u/6

hartnn (hartnn):

not, e^(1/x) /6

OpenStudy (amistre64):

i mean ....\[\int_{a}^{b}f(x)~dx=F(b)-F(a)\]

hartnn (hartnn):

plus you need to change limits beforehand, u=1/x when x=1,u=1 when x=2, u=1/2

OpenStudy (anonymous):

but u=1/x?

OpenStudy (amistre64):

if you do a usub, then you need to do hartnns route otherwise its just e^(1/x)/6 applied at 1 and 2

OpenStudy (anonymous):

so i put 1 and 2 in for x?

OpenStudy (amistre64):

3 methods that im aware of; complete usub changes limits to us partial usub where you undo the u back to x and just working thru a normal integration were nothing is changed about.... which is what i did

OpenStudy (anonymous):

\[\frac{ e^u }{ 6 }\]

OpenStudy (amistre64):

\[\int_{1}^2f(x)dx=F(2)-F(1)\] and we determined that \[F(x)=\frac16e^{1/x}\]

OpenStudy (anonymous):

and then i got that^

OpenStudy (anonymous):

so i sub in 2 and 1 for x then subtract them? is that it?

OpenStudy (amistre64):

yes\[\frac16(e^{1/2}-e^{1/1})\] yes\[\frac16(e^{1/2}-e)\]

OpenStudy (anonymous):

if i work that out it gives me something like -0.782?

OpenStudy (anonymous):

if i just leave it as \[\frac{ 1 }{ 6 } (e^\frac{ 1 }{ 2 } - e\]

hartnn (hartnn):

did u mean -0.1782 ? and you can keep your answer as 1/6(e^(1/2)-e)

OpenStudy (anonymous):

when i worked it out it was -0.782 but if i leave it as the 1/6... it works out just the same :)

OpenStudy (anonymous):

thanks for the help :)